Runs with τ_{AD} = 0.01 (black), 0.1 (blue), and 1 (red).

St_{AD} = \frac{n_{n}}{n_{i}} \frac{u_{rms}k_{f}}{ν_{in}}  (1)

Figure 1: Evolution of ξ_{M}(t) and E_{M}(t).

$Header: /var/cvs/brandenb/tex/mhd/ADdecay/notes.tex,v 1.1 2019/12/31 02:34:10 brandenb Exp $